Circular No. 8112

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003df Further to IAUC 8111, W. M. Wood-Vasey, G. Aldering, and P. Nugent

report the discovery of an apparent supernova (mag 20.3) on unfiltered NEAT images taken during Mar. 9.47–9.49 UT. The new object is located at α = 12^{h}07^{m}31^{s}.36, δ = +26^{o}57′36′′.2 (equinox 2000.0), which is 2′′ west and 8′′ north of the center of the apparent host galaxy. Additional magnitudes for SN 2003df from NEAT images: 2002 May 4.15, [21.5; 2003 Mar. 21.45, 20.7; Apr. 8.15, 20.7.

IGR J17464–3213 = XTE J17464–3213 D. Baba and T. Nagata, Nagoya University, on behalf of the Infrared

Survey Facility and Simultaneous Three-color Infrared Imager for Unbiased Survey teams of Nagoya University and the National Astronomical Observatory of Japan; I. Iwata and T. Kato, Kyoto University; and H. Yamaoka, Kyushu University, report on near-infrared (J, H, K_{s}) imaging of the field of the x-ray transient IGR J17464–3213. Two distinct infrared sources exist within 2′′ of the radio counterpart (IAUC 8105); the northeast one (with magnitudes K_{s} = 13.9 ± 0.2 on Apr. 5.97 UT and K_{s} = 13.6 ± 0.2 on Apr. 6.97) is not visible on public 2MASS images, and the southwest one has similar brightness to the source on the 2MASS image. Given the proximity to the radio counterpart and the apparent absence on the 2MASS images, the northeast source (at α = 17^{h}46^{m}15^{s}.57, δ = −32^{o}14′01′′.1, equinox 2000.0; from 34 2MASS reference stars; rms error 0′′.1) is likely to be the infrared counterpart of IGR J17464–3213.

M. P. Rupen, A. J. Mioduszewski, and V. Dhawan, National Radio Astronomy Observatory (NRAO), report a strong radio flare from this xray transient and blackhole candidate (IAUC 8105). Observations with the Very Large Array (VLA) on Apr. 6 and 8 UT show that the source increased in flux density by almost a factor 5 at 4.9 GHz, from 20.1 ± 0.3 to 96 ± 2 mJy. The radio spectrum also changed significantly, from flattish (optically thick) on Apr. 6 to falling with frequency (optically thin) on Apr. 8, with flux densities varying from 19.5 ± 0.3 to 68 ± 2 mJy at 8.5 GHz, 20 ± 1 to 54 ± 2 mJy at 14.9 GHz, 22 ± 2 to 51 ± 2 mJy at 22.5 GHz, and 17 ± 3 to 42 ± 4 mJy at 43.4 GHz. The Apr. 8 spectrum, in particular, is not well fit by a single power law, in the sense that the flux density is too high at the highest frequencies. This suggests the existence of several overlapping components (e.g., an extended jet and a compact core) and may indicate that another flare is on the way.

2003 April 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green